### v2

@@ Abstract
[Abstract]Since the development of the conjugate gradient (CG) method in 1952 by Hestenes and Stiefel, CG, has become an indispensable tool in computational mathematics for solving positive definite linear systems. On the other hand, the conjugate residual (CR) method, closely related CG and introduced by Stiefel in 1955 for the same settings, remains relatively less known outside the numerical linear algebra community. Since their inception, these methods – henceforth collectively referred to as conjugate direction methods – have been extended beyond positive definite to indefinite, albeit consistent, settings. Going one step further, in this paper, we investigate the theoretical and empirical properties of these methods under inconsistent systems. Among other things, we show that small modifications to the original algorithms allow for the pseudo-inverse solution. Furthermore, we show that CR is essentially equivalent to the minimum residual method, proposed by Paige and Saunders in 1975, in such contexts. Lastly, we conduct a series of numerical experiments to shed lights on their numerical stability (or lack thereof) and their performance for inconsistent systems. Surprisingly, we will demonstrate that, unlike CR and contrary to popular belief, CG can exhibit significant numerical instability, bordering on catastrophe in some instances.
keywords:
conjugate direction methods, conjugate gradient, conjugate residual, minimum residual, inconsistent systems, indefinite matrix, pseudo-inverse solution
articletype: Research Articlejournal: Journalvolume: 00
@@ 1 Introduction
where $\mathbf{A}\in\mathbb{R}^{d\times d}$ is a symmetric matrix and $\mathbf{b}\in\mathbb{R}^{d}$. In 1952, Hestenes and Stiefel 1 developed the celebrated conjugate gradient (CG) method to solve 1 when $\mathbf{A}$ is positive definite (PD). Since then, CG has become a cornerstone and essential tool in many areas of applied mathematics, statistics, machine learning, and numerical optimization. For example, in solving for numerical solutions of partial differential equations, the discretization of a differential operator often leads to a positive definite matrix, making CG the go-to method. In statistical modeling and machine learning, CG is frequently the primary solver in methods involving kernel regression. In numerical optimization, the classical Newton-CG method 2 has been the canonical example of truncated Newton-type methods for large-scale problems. A closely related technique to CG is the conjugate residual (CR) method, which was developed by Stiefel in 1955 3 for the same PD settings [^1] In fact, without explicitly mentioning the name, the structure of the method is developed in the original paper of Hestenes and Stiefel 1.. However, despite its profound connections to CG, the CR method remains relatively less known outside the numerical linear algebra community. For example, to our knowledge, the only attempt at employing CR within Newton-type optimization methods is Dahito et al 4.
@@
Both these methods, henceforth collectively referred to as conjugate direction (CD) methods [^2] Precisely speaking, the conjugate direction approach serves as a framework for deriving various methods, including CG and CR. However, in this paper, when we refer to CD, we specifically mean these two particular instances., are examples of Krylov subspace algorithms for solving 1 when $\mathbf{A}$ is PD. However, advancements in science and engineering have led to a growing need for iterative solvers that can effectively address a broader range of challenges beyond PD systems. This includes scenarios where the matrix $\mathbf{A}$ is symmetric, but singular and/or indefinite, as well as inconsistent settings, namely $\mathbf{b}\not\in\text{Range}(\mathbf{A})$, e.g., 5, 6, 7, 8. As a result, the popularity of CD, especially CG, to solve problems under these conditions continues to extend. There are studies 9, 10, 11, 12, 13, 14 providing comprehensive treatment of CD beyond positive definite systems. However, all these papers operate under the assumption that the system is consistent, i.e. $\mathbf{b}\in\text{Range}(\mathbf{A})$.
@@
In contrast, the minimum residual [^3] We would like to caution the reader that Algorithm 5.3 in the book 15 is called Minimal Residual Iteration which is entirely unrelated to MINRES. (MINRES) method by Paige and Saunders 16 extends naturally to all systems involving symmetric matrices, without relying on a consistency assumption and has shown great success in numerical optimization in recent years, e.g., 7, 17, 18, 19. In comparison to CG or CR, however, MINRES is a significantly more complex algorithm and entails more intricate implementations. Therefore, understanding the properties and the behaviors of the CD methods in inconsistent settings can be of great interest. Before delving any deeper, it is essential to note that our objective in this paper is not to devise novel variants of CG or CR that can extend to inconsistent systems. Instead, our emphasis is on exploring the theoretical and empirical behavior of the original methods in such situations.
@@
    We first show that the classical CG is fundamentally unable to recover any solution of the normal equation (Theorem 3.1). Subsequently, we provide a modification (Algorithm 3) that ensures convergence to the pseudo-inverse solution (Theorem 3.8).
@@
    We provide a variant of CR (Algorithm 4) that converges to the pseudo-inverse solution (Theorem 3.18). We also establish that, in essence, CR and MINRES are equivalent (Theorem 3.23).
@@
    We numerically demonstrate that, unlike CR and contrary to popular belief, CG can be numerically unstable in a borderline catastrophe manner, in many instances (Section 4).
@@
(Lemma 3.3) (Lemma 3.11)
Table 1: Similarities and differences between CG (Algorithm 1) and CR (Algorithm 2) at termination. Here, $\mathbf{x}$ is the final iterate, $\mathbf{r}=\mathbf{b}-\mathbf{Ax}$ is the residual at $\mathbf{x}$, $\mathbf{p}$ is the update direction within the CD methods, $\mathbf{x}^{+}$ and $\mathbf{x}^{\star}$ denote the pseudo-inverse solution and a normal solution with $\mathbf{y}\neq\mathbf{0}$ from Definition 1.1, respectively. All quantities are at the termination of the respective algorithm.
The rest of this paper is organized as follows. We end this section by introducing the notation and definitions that are used throughout this work. We first review the CD methods for consistent systems in Section 2. Subsequently, we shift our focus to the central theme of our paper and present theoretical results for inconsistent systems in Section 3. Finally, in Section 4, we provide numerical experiments to verify our theoretical results as well as to shed light on the numerical performance of CD methods in various situations.
@@
Definition 1.1 (Normal Solutions and the Pseudo-inverse Solution).
@@
The Krylov subspace of degree $k\geq 1$, generated by $\mathbf{A}$ and $\mathbf{v}$, is denoted as $\mathcal{K}_{k}(\mathbf{A},\mathbf{v})=\text{Span}\{\mathbf{v},\mathbf{Av},\cdots\mathbf{A}^{k-1}\mathbf{v}\}$. Recall that the grade of $\mathbf{v}$ with respect to $\mathbf{A}$ is defined as follows 20.
Definition 1.2 (Grade of $\mathbf{v}$ with respect to $\mathbf{A}$).
@@
Remark 1.4 (Symmetric Matrix and Non-zero Right-hand side).
@@
Definition 1.5 (Zero-curvature Direction).
@@ 2 Consistent Systems: Review
Equivalently, one can derive $\mathbf{x}_{k}$ by enforcing the Petrov-Galerkin condition 20, as the vector in the search space $\mathcal{K}_{k}(\mathbf{A},\mathbf{b})$ whose residual $\mathbf{r}_{k}$ is orthogonal to a constraint space, which in the case of CG is taken to be the same as $\mathcal{K}_{k}(\mathbf{A},\mathbf{b})$. When $k=g$, i.e., the grade of $\mathbf{b}$ with respect to $\mathbf{A}$, the solution that minimizes 3 is also the solution to 1.
@@
highlighting the well-known equivalence between CR and MINRES for when $\mathbf{A}\succ\mathbf{0}$ 21. We remind that while both CG and CR methods involve one matrix-vector product in each iteration, CR requires one more vector of storage and one more vector update compared with CG. Indeed, in Algorithm 2, after computing $\mathbf{Ar}_{k}$, one can calculate $\mathbf{Ap}_{k}$ as $\mathbf{Ap}_{k}=\mathbf{Ar}_{k}+\beta_{k}\mathbf{Ap}_{k-1}$; see 20 for more details.
@@
Going beyond PD settings, there have been several efforts in extending the CD methods. However, in all these attempts, the underlying assumption is that the linear system is consistent. The extension to consistent systems involving positive semi-definite (PSD) matrices is relatively straightforward and has been explored in the literature 22, 23. In the work of Kaasschieter 24, the behavior of preconditioned CG under positive semi-definite, but consistent, systems is studied. The main challenge in extending the CD methods to indefinite settings is the risk of encountering division by zero. In CG, for example, this occurs when $\left\langle\mathbf{p}_{k-1},\mathbf{Ap}_{k-1}\right\rangle=0$ at which point $\alpha_{k-1}$ is no longer well-defined and CG breaks down. This can happen when either $\mathbf{Ap}_{k-1}=\mathbf{0}$ or $\mathbf{Ap}_{k-1}\perp\mathbf{p}_{k-1}$. We will show later that the former case constitutes a breakdown, as it signals the final iteration and CG can be safely terminated (Lemma 3.3). The latter case, however, is considered an “unlucky breakdown”. For CG applied to indefinite, but consistent, settings, this situation is handled by employing the concept of a hyperbolic pair 9, 10. Likewise, there are works 12, 13, 14, which developed planar CG, offering an alternative approach for handling unlucky breakdowns within CG for indefinite, but consistent, settings. In indefinite settings, similar breakdowns can occur in CR when $\left\langle\mathbf{r}_{k-1},\mathbf{Ar}_{k-1}\right\rangle=0$ at which point $\beta_{k-1}$ is not well-defined and CR breaks down. This can happen when either $\mathbf{Ar}_{k-1}=\mathbf{0}$ or $\mathbf{Ar}_{k-1}\perp\mathbf{r}_{k-1}$. Again, we will later show that the former case is a breakdown that is theoretically guaranteed by Theorem 3.13, while the latter scenario is again characterized as an unlucky breakdown.
@@
As mentioned earlier, our goal in this paper is not to design new variants of CG or CR that can extend to indefinite systems. Instead, our emphasis is on exploring the behavior of the original methods in such situations. As a result, to exclude cases where the algorithm encounters an unlucky breakdown in indefinite settings, we make an umbrella assumption that a zero-curvature, as in Definition 1.5, is not encountered in the algorithm prior to termination. {assumption}[No Unlucky Breakdown] If $\mathbf{A}$ is indefinite, the CD methods do not encounter a zero-curvature direction as in Definition 1.5. We emphasize that Algorithm 2 is only required for indefinite settings. This is because it turns out that for positive semi-definite matrices, the only breakdown is a lucky one, which only happens at the very last iteration, and signals that the algorithm has indeed reached the grade; see Lemmas 3.3 and 3.13. We note in passing that Algorithm 2 is not overly restrictive since with a suitably random starting point, it is highly likely to hold for all practical purposes.
Remark 2.1.
@@
One might suspect that with $\mathbf{x}_{0}$ if, at the first iteration, $\langle\mathbf{b},\mathbf{Ab}\rangle\neq 0$, subsequent iterations would also avoid an unlucky breakdown. However, as noted in 25, this is not necessarily the case. The following example,
@@
Let $\mathbf{b}\in\text{Range}(\mathbf{A})$. Under Algorithm 2, Algorithms 1 and 2 converge to the pseudo-inverse solution $\mathbf{x}_{g}=\mathbf{A}^{\dagger}\mathbf{b}$ and $\mathbf{r}_{g}=\mathbf{0}$.
@@
Let $\mathbf{b}\in\text{Range}(\mathbf{A})$. Under Algorithm 2, for Algorithms 1 and 2, we have $\mathbf{p}_{k}=\mathbf{0}$ if and only if $k=g$.
For consistent systems, under Algorithm 2, both the residual vector (Fact 1), and the update direction (Fact 2) are zero at termination for the CD methods. Fact 1 has been shown in the works of Hayami 22, 23 for semi-definite matrices. It is a straightforward extension to the indefinite matrices with Algorithm 2 in place, and the proof of Fact 1 is a direct consequence of Lemma B.3, which can be found in Appendix B. Fact 2 indicates that, under consistent systems, not only does $\mathbf{r}_{g}=\mathbf{0}$ at termination, but $\mathbf{p}_{g}=\mathbf{0}$ as well. While this fact may seem trivial, it is worth noting as it sharply contrasts with the behavior of CG under inconsistent systems, where $\mathbf{Ar}$ is never zero (Theorem 3.1), but $\mathbf{Ap}$ is zero at termination (Lemma 3.3).

### v4

@@ Abstract
Since the development of the conjugate gradient (CG) method in 1952 by Hestenes and Stiefel, CG, has become an indispensable tool in computational mathematics for solving positive definite linear systems. On the other hand, the conjugate residual (CR) method, closely related CG and introduced by Stiefel in 1955 for the same settings, remains relatively less known outside the numerical linear algebra community. Since their inception, these methods – henceforth collectively referred to as conjugate direction methods – have been extended beyond positive definite to indefinite, albeit consistent, settings. Going one step further, in this paper, we investigate theoretical and empirical properties of these methods under inconsistent systems. Among other things, we show that small modifications to the original algorithms allow for the pseudo-inverse solution. Furthermore, we show that CR is essentially equivalent to the minimum residual method, proposed by Paige and Saunders in 1975, in such contexts. Lastly, we conduct a series of numerical experiments to shed lights on their numerical stability (or lack thereof) and their performance for inconsistent systems. Surprisingly, we will demonstrate that, unlike CR and contrary to popular belief, CG can exhibit significant numerical instability, bordering on catastrophe in some instances.
@@ 1 Introduction
where $\mathbf{A}\in\mathbb{R}^{d\times d}$ is a symmetric matrix and $\mathbf{b}\in\mathbb{R}^{d}$. In 1952, Hestenes and Stiefel [22] developed the celebrated conjugate gradient (CG) method to solve 1 when $\mathbf{A}$ is positive definite (PD). Since then, CG has become a cornerstone and essential tool in many areas of applied mathematics, statistics, machine learning, and numerical optimization. For example, in solving for numerical solutions of partial differential equations, the discretization of a differential operator often leads to a positive definite matrix, making CG the go-to method. In statistical modeling and machine learning, CG is frequently the primary solver in methods involving kernel regression. In numerical optimization, the classical Newton-CG method [33] has been the canonical example of truncated Newton-type methods for large-scale problems. A closely related technique to CG is the conjugate residual (CR) method, which was developed by Stiefel in 1955 [43] for the same PD settings [^4] In fact, without explicitly mentioning the name, the structure of the method is developed in the original paper of Hestenes and Stiefel [22].. However, despite its profound connections to CG, the CR method remains relatively less known outside the numerical linear algebra community. For example, to our knowledge, the only attempt at employing CR within Newton-type optimization methods is Dahito et al [6].
@@
Both these methods, henceforth collectively referred to as conjugate direction (CD) methods [^5] Precisely speaking, the conjugate direction approach serves as a framework for deriving various methods, including CG and CR. However, in this paper, when we refer to CD, we specifically mean these two particular instances., are examples of Krylov subspace algorithms for solving 1 when $\mathbf{A}$ is PD. However, advancements in science and engineering have led to a growing need for iterative solvers that can effectively address a broader range of challenges beyond PD systems. This includes scenarios where the matrix $\mathbf{A}$ is symmetric, but singular and/or indefinite, as well as inconsistent settings, namely $\mathbf{b}\not\in\text{Range}(\mathbf{A})$, e.g., [44, 38, 36, 31]. As a result, the popularity of CD, especially CG, to solve problems under these conditions continues to extend. There are studies [12, 29, 30, 8, 9, 10, 15, 42] providing comprehensive treatment of CD beyond positive definite systems. However, all these results operate under the assumption that the system is consistent, i.e. $\mathbf{b}\in\text{Range}(\mathbf{A})$.
@@
In contrast, the minimum residual [^6] We would like to caution the reader that Algorithm 5.3 in the book [40] is called Minimal Residual Iteration which is entirely unrelated to MINRES. (MINRES) method by Paige and Saunders [34] extends naturally to all systems involving symmetric matrices, without relying on a consistency assumption and has shown great success in numerical optimization in recent years, e.g., [36, 24, 27, 28]. In comparison to CG or CR, however, MINRES is a significantly more complex algorithm and entails more intricate implementations. Therefore, understanding the properties and the behavior of the CD methods in inconsistent settings can be of great interest. Before delving any deeper, it is essential to note that our objective in this paper is not to devise novel variants of CG or CR that can extend to inconsistent systems. Instead, our emphasis is on exploring the theoretical and empirical behavior of the original methods in such situations.
@@
    We first show that the classical CG is fundamentally unable to recover any solution of the normal equation (Theorem 1). Subsequently, we provide a modification (Algorithm 3) that ensures convergence to the pseudo-inverse solution (Theorem 2).
@@
    We provide a variant of CR (Algorithm 4) that converges to the pseudo-inverse solution (Theorem 4). We also establish that, in essence, CR and MINRES are equivalent (Theorem 5).
@@
    We numerically demonstrate that, beyond typical finite precision errors, CG–unlike CR–can exhibit borderline catastrophic instability in inconsistent settings. These instabilities can persist even for PD matrices with a few small eigenvalues.
@@
(Lemma 1) (Lemma 3)
Table 1: Similarities and differences between CG (Algorithm 1) and CR (Algorithm 2) at termination. Here, $\mathbf{x}$ is the final iterate, $\mathbf{r}=\mathbf{b}-\mathbf{Ax}$ is the residual at $\mathbf{x}$, $\mathbf{p}$ is the update direction within the CD methods, $\mathbf{x}^{+}$ and $\mathbf{x}^{\star}$ denote the pseudo-inverse solution and a normal solution with $\mathbf{y}\neq\mathbf{0}$ from Definition 1, respectively. All quantities are at the termination of the respective algorithm.
The rest of this paper is organized as follows. We end this section by introducing the notation and definitions that are used throughout this work. We first review the CD methods for consistent systems in Section 2. Subsequently, we shift our focus to the central theme of our paper and present theoretical results for inconsistent systems in Section 3.2.2. Finally, in Section 4, we provide numerical experiments to verify our theoretical results as well as to shed light on the numerical performance of CD methods in various situations.
@@
Definition 1 (Normal Solutions and the Pseudo-inverse Solution).
@@
The Krylov subspace of degree $k\geq 1$, generated by $\mathbf{A}$ and $\mathbf{v}$, is denoted as
$\mathcal{K}_{k}(\mathbf{A},\mathbf{v})=\text{Span}\{\mathbf{v},\mathbf{Av},\cdots,\mathbf{A}^{k-2}\mathbf{v},\mathbf{A}^{k-1}\mathbf{v}\}.$
Recall that the grade of $\mathbf{v}$ with respect to $\mathbf{A}$ is defined as follows [1].
Definition 2 (Grade of $\mathbf{v}$ with respect to $\mathbf{A}$).
@@
Remark 2 (Symmetric Matrix and Non-zero Right-hand side).
@@
Definition 3 (Zero-curvature Direction).
@@ 2 Consistent Systems: Review
Equivalently, one can derive $\mathbf{x}_{k}$ by enforcing the Petrov-Galerkin condition [1], as the vector in the search space $\mathcal{K}_{k}(\mathbf{A},\mathbf{b})$ whose residual $\mathbf{r}_{k}$ is orthogonal to a constraint space, which in the case of CG is taken to be the same as $\mathcal{K}_{k}(\mathbf{A},\mathbf{b})$. When $k=g$, i.e., the grade of $\mathbf{b}$ with respect to $\mathbf{A}$, the solution that minimizes 3 is also the solution to 1.
@@
highlighting the well-known equivalence between CR and MINRES for when $\mathbf{A}\succ\mathbf{0}$ [14]. We remind that while both CG and CR methods involve one matrix-vector product in each iteration, CR requires one more vector of storage and one more vector update compared with CG. Indeed, in Algorithm 2, after computing $\mathbf{Ar}_{k}$, one can calculate $\mathbf{Ap}_{k}$ as $\mathbf{Ap}_{k}=\mathbf{Ar}_{k}+\beta_{k}\mathbf{Ap}_{k-1}$; see [1] for more details.
@@
Going beyond PD settings, there have been several efforts in extending the CD methods. However, in all these attempts, the underlying assumption is that the linear system is consistent. The extension to consistent systems involving positive semi-definite (PSD) matrices is relatively straightforward and has been explored in the literature [15, 19, 18]. In the work of Kaasschieter [23], the behavior of preconditioned CG under positive semi-definite, but consistent, systems is studied. The main challenge in extending the CD methods to indefinite settings is the risk of encountering division by zero. In CG, for example, this occurs when $\left\langle\mathbf{p}_{k-1},\mathbf{Ap}_{k-1}\right\rangle=0$ at which point $\alpha_{k-1}$ is no longer well-defined and CG breaks down. This can happen when either $\mathbf{Ap}_{k-1}=\mathbf{0}$ or $\mathbf{Ap}_{k-1}\perp\mathbf{p}_{k-1}$. We will show later that the former case constitutes a breakdown, as it signals the final iteration and CG can be safely terminated (Lemma 1). The latter case, however, is considered an “unlucky breakdown”. For CG applied to indefinite, but consistent, settings, this situation is handled by employing the concept of a hyperbolic pair [12, 29]. Likewise, there are works [8, 9, 10], which developed planar CG, offering an alternative approach for handling unlucky breakdowns within CG for indefinite, but consistent, settings. In indefinite settings, similar breakdowns can occur in CR when $\left\langle\mathbf{r}_{k-1},\mathbf{Ar}_{k-1}\right\rangle=0$ at which point $\beta_{k-1}$ is not well-defined and CR breaks down. This can happen when either $\mathbf{Ar}_{k-1}=\mathbf{0}$ or $\mathbf{Ar}_{k-1}\perp\mathbf{r}_{k-1}$. Again, we will later show that the former case is a breakdown that is theoretically guaranteed by Theorem 3, while the latter scenario is again characterized as an unlucky breakdown.
@@
As mentioned earlier, our goal in this paper is not to design new variants of CG or CR that can extend to indefinite systems. Instead, our emphasis is on exploring the convergent behavior of the original methods in such situations. As a result, to exclude cases where the algorithm encounters an unlucky breakdown in indefinite settings, we make an umbrella assumption that a zero-curvature, as in Definition 3, is not encountered in the algorithm prior to termination.
Assumption 1 (No Unlucky Breakdown).
If $\mathbf{A}$ is indefinite, the CD methods do not encounter a zero-curvature direction as in Definition 3.
Without Assumption 1, the convergence behavior of CG and CR for inconsistent systems can generally be studied based on whether or not an unlucky breakdown occurs during the iterations. In the case of an unlucky breakdown, both CG and CR fail to converge by definition, which is an uninteresting scenario. The more compelling question is to explore their convergence behavior in the absence of an unlucky breakdown throughout the iterations. Therefore, to align with our goal of investigating the convergence of CG and CR for inconsistent indefinite systems, we exclude the former case and focus on the latter, justifying the introduction of Assumption 1. Furthermore, we note that Assumption 1 is only required for indefinite settings. This is because it turns out that for positive semi-definite matrices, the only breakdown is a lucky one, which only happens at the very last iteration, and signals that the algorithm has indeed reached the grade; see Lemmas 1 and 3. In addition, in most practical applications, unlucky breakdowns are virtually never encountered, which is consistent with the observations in [42]. Therefore, we argue that Assumption 1 is reasonable in our context.
We note in passing that, in some applications, encountering a zero-curvature direction while solving 1 can be algorithmically beneficial. For instance, in Newton-type optimization methods, where $\mathbf{A}$ and $\mathbf{b}$ typically represent the Hessian and the negative gradient of a function, respectively, a zero-curvature direction can be leveraged to construct effective descent directions (see, e.g., [27, 24, 39]).
Remark 3.
@@
One might suspect that with $\mathbf{x}_{0}$ if, at the first iteration, $\langle\mathbf{b},\mathbf{Ab}\rangle\neq 0$, subsequent iterations would also avoid an unlucky breakdown. However, as noted in [13], this is not necessarily the case. The following example,
@@
Let $\mathbf{b}\in\text{Range}(\mathbf{A})$. Under Assumption 1, Algorithms 1 and 2 converge to the pseudo-inverse solution $\mathbf{x}_{g}=\mathbf{A}^{\dagger}\mathbf{b}$ and $\mathbf{r}_{g}=\mathbf{0}$.
@@
Let $\mathbf{b}\in\text{Range}(\mathbf{A})$. Under Assumption 1, for Algorithms 1 and 2, we have $\mathbf{p}_{k}=\mathbf{0}$ if and only if $k=g$.
For consistent systems, under Assumption 1, both the residual vector (Fact 1), and the update direction (Fact 2) are zero at termination for the CD methods. Fact 1 has been shown in the works of Hayami [19, 18] for semi-definite matrices. It is a straightforward extension to the indefinite matrices with Assumption 1 in place, and the proof of Fact 1 is a direct consequence of Lemma 15, which can be found in Appendix B. Fact 2 indicates that, under consistent systems, not only does $\mathbf{r}_{g}=\mathbf{0}$ at termination, but $\mathbf{p}_{g}=\mathbf{0}$ as well. While this fact may seem trivial, it is worth noting as it sharply contrasts with the behavior of CG under inconsistent systems, where $\mathbf{Ar}$ is never zero (Theorem 1), but $\mathbf{Ap}$ is zero at termination (Lemma 1).